OpenStudy (anonymous):

Identify the oblique asymptote of f(x) = quantity x plus 4 over quantity 3 x squared plus 5 x minus 2.

Nnesha (nnesha):

if the degree of the numerator is one less than the degree of the denominator then you can find oblique asymptote

Nnesha (nnesha):

\[f(x) =\frac{x+4}{3x^2+5x-2}\] like this ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

alright then read my first comment slant/oblique asy occurs when the highest degree of the denominator is one less than the numerator

Nnesha (nnesha):

ignore my first comment i messed up /.\

OpenStudy (anonymous):

so it it not an oblique

OpenStudy (anonymous):

wait im still confused

Nnesha (nnesha):

alright why ?

OpenStudy (anonymous):

because you said for it to be an oblique its degree of the denominator is one less than the numerator

Nnesha (nnesha):

slant/oblique asy occurs when the highest degree of the denominator is one less than the numerator here is an example \[\huge\rm \frac{ 2x^1+1 }{ x^2 +3x+4 }\] higest degree of the denominator is greater than the numerator so then there is no oblique asy

OpenStudy (anonymous):

thats what i said

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Motown117 so it it not an oblique \(\color{blue}{\text{End of Quote}}\) f(x) doesn't hve oblique asy yes let's say if the equation \[f(x)= \frac{ 3x^2+5x+2 }{ x+4 }\]then you can use either long or synthetic division to find slant/oblique asy y = quotient< == oblique asy

OpenStudy (anonymous):

but noone responed to my answer

Nnesha (nnesha):

i was typing at the that moment too much lag it's hard to use equation tool &ect

OpenStudy (anonymous):

oh srry but thank you for the help tho

Nnesha (nnesha):

np :=)) glad to hlp

OpenStudy (anonymous):

wait what?

Nnesha (nnesha):

oblique asy not horizontal

Nnesha (nnesha):

oblique or slant is different than horizontal asy :=))

OpenStudy (anonymous):

@Nnesha can you help me with vertical asymptotes

Nnesha (nnesha):

sure

OpenStudy (anonymous):

Identify the vertical asymptotes of f(x) = quantity x plus 6 over quantity x squared minus 9x plus 18. x = −6 and x = −3 x = −6 and x = 3 x = 6 and x = −3 x = 6 and x = 3

Nnesha (nnesha):

for \(\color{green}{\rm Vertical~ asy.}\) set the denominator equal to zero and then solve for the variable.

OpenStudy (anonymous):

so x=6 and x=3

Nnesha (nnesha):

\[\frac{ x+6 }{ x^2+9x+18 }\] like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^2-9x+18

Nnesha (nnesha):

alright then there is a quadratic equation first factor that

Nnesha (nnesha):

x^2+9x+18 AC method is easy to factor it what two numbers would you multiply to get C which is 18 and when u add them u should get middle term b which is 9

OpenStudy (anonymous):

6 times 3

Nnesha (nnesha):

right and since the lading coefficient is one so (x+ 6)(x+3 are factors \[\frac{ x+6 }{ (x+6)(x+3) }\] first simplify and then set the denomiantor equal to zero

OpenStudy (anonymous):

so its D then right

Nnesha (nnesha):

hmm weird there should be only one x value

Nnesha (nnesha):

i think they don't want us to factor the quadratic equation x^2+9x+18=0 so (x+6)(x+3)=0 set each parentheses equal to zero then solve for x

OpenStudy (anonymous):

x=-6 x=-3

Nnesha (nnesha):

looks right but if we the other way \[\frac{ \cancel{x+6 }}{ \cancel{(x+6)}(x+3) }\] cancel the (x+6) and then x+3=0 would be asy hmm

OpenStudy (anonymous):

thank you again

Nnesha (nnesha):

yw